A box contains a certain no. of balls, on each of 60% balls A is marked and on each of 30% balls letter B is marked and on each of the remaining balls letter C idols marked . A ball is drawn at random, find the probability that the ball drawn is (1)marked C
(2)marked A or B
(3)marked neither B nor C
1
Expert's answer
2019-07-16T09:57:03-0400

(1) because the probability to get A ball 60% and B 30% C=100-30-60=10% => P(C)=0.1;

(2) P(A+B)=P(A)+P(B)=0.6+0.3=0.9;

(3) because all balls are marked then balls are not B and not C will be A=> P=P(A)=0.6.